OpenStudy (petlovers123):

What is the remainder when 15,908 is divided by 13? @Zepdrix

zepdrix (zepdrix):

Long division? :)

OpenStudy (petlovers123):

thx

hartnn (hartnn):

would the remainder with 13 be same as \(15908 -13000 = 2908\) it would be much easier to calculate, right?

hartnn (hartnn):

remainder of (15908) - remainder of 13000 = remainder of (15908) -0 = remainder of (15908) . similarly 2908 - 2600 = 308

hartnn (hartnn):

308 - 260 = 48 now whats the remainder when 48 is divided by 13?

hartnn (hartnn):

or go one more step :P 48 - 39 = 9 and done :P

OpenStudy (petlovers123):

Thank you so much 

OpenStudy (petlovers123):

I became your 2800

OpenStudy (petlovers123):

fan

hartnn (hartnn):

awesome! thanks :D

hartnn (hartnn):

I became your 10th! :)